OpenStudy (anonymous):

Rewrite the rational exponents:^6√xy^5z

jhonyy9 (jhonyy9):

missed the first term

jimthompson5910 (jim_thompson5910):

Remember that \[\large \sqrt[n]{x}=x^{\frac{1}{n}}\] Using this, we can say that \[\large \sqrt[6]{xy^5z}=\left(xy^5z\right)^{\frac{1}{6}}\]